# Improved Extended Regular Expression Matching

### v1

@@ Abstract
An extended regular expression $R$ specifies a set of strings formed by characters from an alphabet combined with concatenation, union, intersection, complement, and star operators. Given an extended regular expression $R$ and a string $Q$, the extended regular expression matching problem is to decide if $Q$ matches any of the strings specified by $R$. Extended regular expressions are a basic concept in formal language theory and a basic primitive for searching and processing data. Extended regular expression matching was introduced by Hopcroft and Ullmann in the 1970s [Introduction to Automata Theory, Languages and Computation, Addison-Wesley, 1979], who gave a simple dynamic programming solution using $O(n^{3}m)$ time and $O(n^{2}m)$ space, where $n$ is the length of $Q$ and $m$ is the length of $R$. Since then, several solutions have been proposed, but few significant asymptotic improvements have been obtained. The current state-of-the art solution, by Yamamoto and Miyazaki [COCOON, 2003], uses $O(\frac{n^{3}k+n^{2}m}{w}+n+m)$ time and $O(\frac{n^{2}k+nm}{w}+n+m)$ space, where $k$ is the number of negation and complement operators in $R$ and $w$ is the number of bits in a word. This roughly replaces the $m$ factor with $k$ in the dominant terms of both the space and time bounds of the classic Hopcroft and Ullmann algorithm.
@@
$O\left(n^{\omega}k+\frac{n^{2}m}{\min(w/\log w,\log n)}+m\right)$
time and $O(\frac{n^{2}\log k}{w}+n+m)=O(n^{2}+m)$ space, where $\omega\approx 2.3716$ is the exponent of matrix multiplication. Essentially, this replaces the dominant $n^{3}k$ term with $n^{\omega}k$ in the time bound, while simultaneously improving the $n^{2}k$ term in the space to $O(n^{2})$.
To achieve our result, we develop several new insights and techniques of independent interests, including a new compact representation to store and efficiently combine substring matches, a new clustering technique for parse trees of extended regular expressions, and a new efficient combination of finite automaton simulation with substring match representation to speed up the classic dynamic programming solution.
@@ 1 Introduction
An extended regular expression $R$ specifies a set of strings formed by characters from an alphabet combined with concatenation ($\odot$), union ($\mid$), intersection ($\cap$), complement ($\neg$), and star () operators. For instance, $\neg((a|b)^{*})\odot b$ describes the set of strings that end in $b$ and contains at least one character that is not $a$ or $b$. Given an extended regular expression $R$ and a string $Q$, the extended regular expression matching problem is to decide if $Q$ matches any of the strings specified by $R$. The intersection and complement are the extended operators, and without these, the expression is simply called a regular expression. Regular expressions are a fundamental concept in formal language theory introduced by Kleene in the 1950s [30], and extended and non-extended regular expression matching is a basic tool in computer science for searching and processing text. Standard tools such as grep and sed provide direct support for extended regular expression matching in files, and the scripting language perl [52] is a full programming language designed to support extended regular expression matching easily. Extended regular expression matching appears in many large-scale data processing applications such as internet traffic analysis [28, 58, 32, 23, 36, 51, 11, 2, 5], data mining [21, 1, 50, 15], databases [34, 38, 26, 37, 18], computational biology [40, 41, 3, 54, 17], and human-computer interaction [29, 47, 48].
@@
The extended regular expression matching problem has also attracted significant research interest [25, 24, 31, 33, 56, 55, 27, 57, 44, 42, 43], but in contrast to non-extended regular expression matching, surprisingly little is known about the complexity of the problem. The problem was first introduced in 1979 by Hopcroft and Ullmann [25] as a textbook exercise (see exercise 3.23 and the associated solution). They presented a simple dynamic programming algorithm using $O(n^{3}m)$ time and $O(n^{2}m)$ space. However, with careful implementation and modern matrix multiplication algorithms, it is straightforward to improve the bound to $O(n^{\omega}m)$ time and $O(n^{2}m/w+n+m)$ space. Here, $\omega\approx 2.3716$ is the exponent of matrix multiplication [53] and $w$ is the length of a machine word. Apparently, these improvements have been overlooked by later papers. We review this algorithm in detail and show how to implement it efficiently in Section 3.
@@
Several improved algorithms have been proposed [24, 31, 33, 56, 55, 27, 57, 44, 42, 43]. Hirst in 1989 [24], and later Kupferman and Zuhovitzky in 2002 [33], and Ilie, Shan, and Yu in 2003 [27] all proposed solutions using $O(n^{2}m)$ time. Unfortunately, these results were all found to contain errors, and their claimed complexity bounds were incorrect. See Roşu [43, 42] for a detailed discussion of this.
@@
Only a few improvements of the classic dynamic programming solution (without errors) are known. Let $n$ and $m$ be the lengths of the string and the extended regular expression, respectively. Yamamoto and Miyazaki [57] gave an $O(\frac{n^{3}k+n^{2}m}{w}+n+m)$ time and $O(\frac{n^{2}k+nm}{w}+n+m)$ space algorithm, where $k$ is the number of negation and complement operators in the regular expression. Essentially, if we ignore the $w$ factor, this result replaces the $m$ factor with $k$ in the dominant term of both the time and space bounds of the original $O(n^{3}k)$ time and $O(n^{2}m)$ for most combinations of parameters. This result uses techniques different from the original dynamic programming algorithm and it does not appear that the bound can be easily improved using matrix multiplication. Yamamoto [56] also gave an $O(n^{3}\lambda+n^{2}m)$ time and $O(n^{2}\lambda+nm)$ space algorithm, where $\lambda$ is the number of critical modules of the regular expression. In general, a critical module is a subexpression of $R$ that contains a single negation or intersection operator. However, a module is only critical depending on its placement relative to other modules and star operators, and hence $\lambda\leq k$. Thus, compared to Yamamoto and Miyazaki [57] (again ignoring the $w$ factor), the $k$ factor is replaced by $\lambda$. As before, it does not appear that this bound can be improved using matrix multiplication. Finally, Roşu [43] proposed an incomparable result that uses exponential time and space in $m$.
1.1 Results
@@
$O\left(n^{\omega}k+\frac{n^{2}m}{\min(w/\log w,\log n)}+m\right).$
Theorem 1 improves the time and space bounds of all the previous results for almost all combinations of parameters. For the time bound, we either replace the dependency on $m$ with $k$ compared to the dynamic programming solution [25] (implemented with fast matrix multiplication) or improve the dominant $n^{3}k/w$ term by a factor $n^{3-\omega}/w\approx n^{0.6284}/w$ compared to Yamamoto and Miyazaki [57]. For the space bound, we obtain the first $O(n^{2})$ space solution, improving the previous results by a factor $m$ or $k/w$. A key component in the dynamic programming algorithm and our algorithm is storing $(n+1)\times(n+1)$ boolean matrices for fast matrix multiplication. Note that this requires $\Omega(n^{2})$ bits space, and we match this bound within a logarithmic factor.
We actually show a more general result than Theorem 1, which takes a black-box implementation of a (non-extended) regular expression matching algorithm (subject to a few natural constraints) and uses it within our general solution to extended regular expression matching. To obtain the bounds in Theorem 1, we plug in efficient algorithms for regular expression matching [8, 7, 13]. Improvements on the non-extended regular expression matching problem thus immediately imply improvements on Theorem 1.
Our focus in this paper are the classic extended operators. However, our framework is sufficiently general to handle other important operators with little overhead. For instance, we can also efficiently support the interval operator $R^{[i,j]}$, which specifies between $i$ and $j$ concatenations of $R$. We will cover these generalizations in detail in an extended version of this paper.
@@ 1.2 Technical Overview
The match graphs allow us to, in a simple way, describe the classic dynamic programming solution of Hopcroft and Ullmann [25] and its improvement. Essentially, the algorithm processes the parse tree of extended regular expression in a bottom-up traversal, computing at each node $v$ a match graph representing the substrings of $Q$ that match the subexpression of $R$ represented by $v$. Using our match graph operations, we explain how to reduce the time from $O(n^{3}m)$ to $O(n^{\omega}m)$.
Next, we show how to cluster the parse tree of $R$ into a hierarchy of connected subtrees, called clusters, such that each leaf cluster contains no node corresponding to an extended operator, and each internal cluster contains exactly one node that either corresponds to an extended operator or is the lowest common ancestor of a pair of extended operators, thus guaranteeing that we have no more than $O(k)$ clusters.
To speed up the classic dynamic programming algorithm, we essentially process the hierarchy of clusters in a bottom-up traversal. At each cluster C, we show how to efficiently compute the match graph representation for the subtree of the parse tree rooted at the root of $C$ using only a constant number of match graph operations and a number of simulations of finite automata of a (non-extended) regular expression corresponding to $C$.
In combination, the above techniques lead to an algorithm for extended regular expression matching using $O(n^{\omega}k+n^{2}m)$ time and $O(n^{2}k/w+n+m)$ space. We show how to improve the space to $O(n^{2}\log k/w+n+m)=O(n^{2}+m)$ by recursively processing the hierarchy of clusters top-down according to the number of descendants of the children of each cluster and carefully discarding match graphs along the way.
Finally, we generalize the finite automaton simulations in our algorithm to work with a black-box implementation of a (non-extended) regular expression matching algorithm. Plugging in current state-of-the-art results then leads Theorem 1.
@@ 1.3 Outline
We review regular expressions and automata in Section 2. We present match graphs and the improved implementation of the Hopcroft and Ullmann classic dynamic programming algorithm in Section 3. Section 4 presents (a simplified version) of our new algorithm, together with our clustering and a generalization of match graphs. Finally, in Section 5 and 6 we show how to reduce the space and implement our algorithm with a black-box finite automaton simulation.
@@ 2 Regular Expressions and Finite Automata
We review the classical concepts used in the paper. For more details, see e.g., Hopcroft and Ullmann [25].
@@
The parse tree $T(R)$ of $R$ (not to be confused with the parse of $Q$ wrt. to $R$) is the rooted, binary tree representing the hierarchical structure of $R$. The leaves of $T(R)$ are labeled by a character from $\Sigma\cup\{\epsilon\}$ and internal nodes are labeled by either $\odot$, $\mid$, $\cap$, $\neg$, or . For simplicity in our presentation, we identify extended regular expressions with their parse tree. We denote the subtree (or equivalently the subexpression) rooted at $v$ by $R(v)$. See Figure 1 for an example. Given two nodes $v$ and $w$ in $T(R)$, the lowest common ancestor, denoted $\mathsf{lca}(v,w)$, is the common ancestor of $v$ and $w$ with the largest depth.
@@
Given a (non-extended) regular expression $R$, we can construct an NFA accepting precisely the strings in $L(R)$ by several classic methods [35, 22, 49]. In particular, Thompson [49] gave the simple, well-known construction shown in Figure 2. We will call an NFA constructed with these rules a Thompson NFA (TNFA). A TNFA $N(R)$ for $R$ has at most $2m$ states, at most $4m$ transitions, and can be computed in $O(m)$ time. We can compute a state-set transition on a single character in $O(m)$ time with a breadth-first search. Hence, it follows that we can solve (non-extended) regular expression matching on $Q$ in $O(nm)$ time and $O(m)$ space using the above state-set transition algorithm.
@@ 3 Dynamic Programming Algorithm
Let $R$ be an extended regular expression of length $m$ and let $Q$ be a string of length $n$. We review Hopcroft and Ullmann’s dynamic programming algorithm [25] and show how to implement it using fast matrix multiplication in $O(n^{\omega}m)$ time and $O(n^{2}m/w+n+m)$ space.
@@ 3.2 Algorithm
Finally, we report that $R$ matches iff the edge $(g_{0},g_{n})$ is in $G(r)$, where $r$ is the root of $T(R)$.
By induction on the parse tree, it follows that the algorithm correctly determines if $R$ matches $Q$.
To implement the algorithm efficiently, we store each match graph $G(v)$ as a boolean $(n+1)\times(n+1)$ matrix, such that entry $(i,j)$ is $1$ iff there is an edge $(g_{i},g_{j})$ in $G(v)$. Consider the cases in the bottom-up traversal. We have that case 2.1 is a single boolean matrix multiplication of the matrices representing $G(v_{1})$ and $G(v_{2})$, and case 3.2 is the transitive closure of the matrix representing $G(v_{1})$, which can also be computed in the same time as a single matrix multiplication [19]. The remaining cases are straightforward to implement in $O(n^{2})$ time. Thus, we use at most $O(n^{\omega})$ time on each node in the parse tree, and hence the total time is $O(n^{\omega}m)$. Since each matrix uses $O(n^{2})$ bits, the total space is $O(n^{2}m/w+n+m)$.

### v2

@@ Abstract
An extended regular expression $R$ specifies a set of strings formed by characters from an alphabet combined with concatenation, union, intersection, complement, and star operators. Given an extended regular expression $R$ and a string $Q$, the extended regular expression matching problem is to decide if $Q$ matches any of the strings specified by $R$. Extended regular expressions are a basic concept in formal language theory and a basic primitive for searching and processing data. Extended regular expression matching was introduced by Hopcroft and Ullman in the 1970s [Introduction to Automata Theory, Languages and Computation, Addison-Wesley, 1979], who gave a simple dynamic programming solution using $O(n^{3}m)$ time and $O(n^{2}m)$ space, where $n$ is the length of $Q$ and $m$ is the length of $R$. Since then, several solutions have been proposed, but few significant asymptotic improvements have been obtained. The current state-of-the art solution, by Yamamoto and Miyazaki [COCOON, 2003], uses $O(\frac{n^{3}k+n^{2}m}{w}+n+m)$ time and $O(\frac{n^{2}k+nm}{w}+n+m)$ space, where $k$ is the number of negation and complement operators in $R$ and $w$ is the number of bits in a machine word. This roughly replaces the $m$ factor with $k$ in the dominant terms of both the space and time bounds of the classical Hopcroft and Ullman algorithm.
@@
$O\left(n^{\omega}k+\frac{n^{2}m}{\max(w/\log w,\log n)}+m\right)$
time and $O(\frac{n^{2}\log k}{w}+n+m)=O(n^{2}+m)$ space, where $\omega\approx 2.3716$ is the exponent of matrix multiplication. Essentially, this replaces the dominant $n^{3}k$ term with $n^{\omega}k$ in the time bound, while simultaneously improving the $n^{2}k$ term in the space to $O(n^{2})$. We also consider the interval operator (also known as the counting operator) and show how to efficiently extend our result to this case.
To achieve our result, we develop several new insights and techniques of independent interest, including a new compact representation to store and efficiently combine substring matches, a new clustering technique for parse trees of extended regular expressions, and a new efficient combination of finite automaton simulation with substring match representation to speed up the classic dynamic programming solution.
@@ 1 Introduction
An extended regular expression $R$ specifies a set of strings formed by characters from an alphabet combined with concatenation ($\odot$), union ($\mid$), intersection ($\cap$), complement ($\neg$), and star () operators. For instance, $\neg((a|b)^{*})\odot b$ describes the set of strings that end in $b$ and contain at least one character that is not $a$ or $b$. Given an extended regular expression $R$ and a string $Q$, the extended regular expression matching problem is to decide if $Q$ matches any of the strings specified by $R$. The intersection and complement are the extended operators, and without these, the expression is simply called a regular expression. Regular expressions are a fundamental concept in formal language theory introduced by Kleene in the 1950s [36], and extended and non-extended regular expression matching is a basic tool in computer science for searching and processing text. Standard tools such as grep and sed provide direct support for extended regular expression matching in files, and the scripting language perl [62] is a full programming language designed to support extended regular expression matching easily. Extended regular expression matching appears in many large-scale data processing applications such as internet traffic analysis [33, 68, 38, 27, 42, 60, 12, 3, 6], data mining [23, 2, 59, 17], databases [40, 45, 31, 43, 20], computational biology [47, 48, 4, 64, 19], and human-computer interaction [35, 56, 57].
@@
The extended regular expression matching problem has also attracted significant research interest [30, 28, 37, 39, 66, 65, 32, 67, 53, 51, 52, 50, 49], but in contrast to non-extended regular expression matching, surprisingly little is known about the complexity of the problem. The problem was first introduced in 1979 by Hopcroft and Ullman [30] as a textbook exercise (see exercise 3.23 and the associated solution). They presented a simple dynamic programming algorithm that runs in $O(n^{3}m)$ time and uses $O(n^{2}m)$ space. However, with careful implementation and modern matrix multiplication algorithms, it is straightforward to improve the bound to $O(n^{\omega}m)$ time and $O(n^{2}m/w+n+m)$ space. Here, $\omega\approx 2.3716$ is the exponent of matrix multiplication [63] and $w$ is the length of a machine word. Apparently, these improvements have been overlooked by later papers. We review this algorithm in detail and show how to implement it efficiently in Section 3.
@@
Several improved algorithms have been proposed [28, 37, 39, 66, 65, 32, 67, 53, 51, 52]. Hirst in 1989 [28], and later Kupferman and Zuhovitzky in 2002 [39], and Ilie, Shan, and Yu in 2003 [32] all proposed solutions that run in $O(n^{2}m)$ time. Unfortunately, these results were all found to contain errors, and their claimed complexity bounds were incorrect. See Roşu [52, 51] for a detailed discussion of this and the conditional lower bound discussed below.
@@
Only a few improvements of the classic dynamic programming solution (without errors) are known. Let $n$ and $m$ be the lengths of the string and the extended regular expression, respectively. Yamamoto and Miyazaki [67] gave an $O(\frac{n^{3}k+n^{2}m}{w}+n+m)$ time and $O(\frac{n^{2}k+nm}{w}+n+m)$ space algorithm, where $k$ is the number of negation and complement operators in the regular expression. Essentially, if we ignore the $w$ factor, this result replaces the $m$ factor with $k$ in the dominant term of both the time and space bounds of the original $O(n^{3}k)$ time and $O(n^{2}m)$ for most combinations of parameters. This result uses techniques different from the original dynamic programming algorithm, and it does not appear that the bound can be easily improved using matrix multiplication. Yamamoto [66] also gave an $O(n^{3}\lambda+n^{2}m)$ time and $O(n^{2}\lambda+nm)$ space algorithm, where $\lambda$ is the number of critical modules of the regular expression. In general, a critical module is a subexpression of $R$ that contains a single negation or intersection operator. However, a module is only critical depending on its placement relative to other modules and star operators, and hence $\lambda\leq k$. Thus, compared to Yamamoto and Miyazaki [67] (again ignoring the $w$ factor), the $k$ factor is replaced by $\lambda$. As before, it does not appear that this bound can be improved using matrix multiplication. Finally, Roşu [52] proposed an incomparable result that uses exponential time and space in $m$.
On the hardness side, Petersen [50] showed that the problem is LOGCFL-complete. Very recently (and after the initial announcement of our result), Nogami and Terauchi [49] showed that there is no $O(n^{\omega-\epsilon}\mathrm{poly}(m))$ time algorithm for any $\epsilon>0$ unless the $k$-clique hypothesis is false ($\mathrm{poly}(m)$ denotes a constant degree polynomial in $m$). Thus, under this hypothesis, we cannot improve the $\omega$ exponent on the dominant $n^{\omega}$ term by any constant. Note that the earlier incorrect $O(n^{2}m)$ time solutions break this lower bound. Nogami and Terauchi also note that a weaker conditional lower bound can be achieved via a recent reduction from $3k$-Clique to context-free language recognition by Abboud et al. [1].
Our main focus is the extended operators (intersection and complement), but we will also consider including the important interval operator (also often called the counting operator), $R^{[a,b]}$, which specifies the set of strings containing between $a$ and $b$ concatenations of $R$. The interval operator is a standard in most of the practical tools and applications mentioned above for extended operation. While the interval operator has been studied from various perspectives, see e.g., [25, 14, 34, 44, 24, 61, 29, 16], we are not aware of any published results on the complexity of extended regular expression matching with the interval operator.
1.1 Main Results
@@
$O\left(n^{\omega}k+\frac{n^{2}m}{\max(w/\log w,\log n)}+m\right).$
Theorem 1 improves the time and space bounds of all the previous results for almost all combinations of parameters. For the time bound, we either replace the dependency on $m$ with $k$ compared to the dynamic programming solution [30] (implemented with fast matrix multiplication) or improve the dominant $n^{3}k/w$ term by a factor $n^{3-\omega}/w\approx n^{0.6284}/w$ compared to Yamamoto and Miyazaki [67]. By the very recent conditional lower bound by Nogami and Terauchi [49], the exponent on our dominant $n^{\omega}$ factor is optimal assuming the $k$-clique hypothesis. For the space bound, we obtain the first $O(n^{2})$ space solution, improving the previous results by a factor $m$ or $k/w$. A key component of both the dynamic programming and our algorithms is storing $(n+1)\times(n+1)$ boolean matrices for fast matrix multiplication. Note that this requires $\Omega(n^{2})$ bits space, and we match this bound within a logarithmic factor.
We actually show a more general result than Theorem 1, which takes a black-box implementation of a (non-extended) regular expression matching algorithm (subject to a few natural constraints) and uses it within our general solution to extended regular expression matching. To obtain the bounds in Theorem 1, we plug in efficient algorithms for regular expression matching [9, 8, 14]. As a consequence, an improvement on the current bounds for the non-extended regular expression matching problem immediately implies improvements on Theorem 1.
Our algorithmic framework is sufficiently general and simple to allow us to easily and efficiently extend it with other operators. To demonstrate this, we show how to efficiently add support for the interval operator. For an extended regular expression $R$ with interval operators, we achieve the same bounds as in Theorem 1 except for an additional factor $\log B$ on the running time for the interval operators, where $B$ is the maximum of the upper bounds in the ranges for interval operators in $R$ (see Theorem 11 for details).
@@ 1.2 Technical Overview
The match graphs allow us to describe, in a simple way, the classic dynamic programming solution of Hopcroft and Ullman [30] and its improvement. Essentially, the algorithm processes the parse tree of extended regular expression in a bottom-up traversal, computing at each node $v$ a match graph representing the substrings of $Q$ that match the subexpression of $R$ represented by $v$. Using our match graph operations, we explain how to reduce the time from $O(n^{3}m)$ to $O(n^{\omega}m)$.
To speed up the dynamic programming algorithm, we combine the match graphs framework with finite automaton simulation. We first decompose the parse tree of $R$ into a hierarchy of connected subtrees, called clusters, such that each leaf cluster contains no node corresponding to an extended operator, and each internal cluster contains exactly one node that either corresponds to an extended operator or is the lowest common ancestor of a pair of extended operators, thus guaranteeing that we have no more than $O(k)$ clusters. We then process the hierarchy of clusters in a bottom-up traversal and compute at each cluster $C$ the match graph representation for the subtree of the parse tree rooted at the root of $C$. To achieve our desired running time the key technical challenge is to efficiently combine the slow match graph computation with fast finite automaton simulation of a (non-extended) regular expression corresponding to $C$. We present a new technique that achieves this using only a constant number of match graph operations and a constant number of finite automata simulations. In combination, the above techniques lead to an algorithm for extended regular expression matching using $O(n^{\omega}k+n^{2}m)$ time and $O(n^{2}k/w+n+m)$ space.
To achieve the full result of Theorem 1 we first improve the space to $O(n^{2}\log k/w+n+m)=O(n^{2}+m)$. To do so we show how recursively processes the hierarchy of clusters top-down according to the number of descendants of the children of each cluster and carefully discarding match graphs along the way. We then generalize the finite automaton simulations in our algorithm to work with a black-box implementation of a (non-extended) regular expression matching algorithm. Plugging in current state-of-the-art results improves the $O(n^{2}m)$ term in the running time by a factor $\max(w/\log w,\log n))$ and thus leads to Theorem 1. Finally, we demonstrate how to extend our result to the interval operator.
@@ 1.3 Outline
We review regular expressions and finite automata in Section 2. We present match graphs and the improved implementation of the Hopcroft and Ullman classic dynamic programming algorithm in Section 3. Section 4 presents (a simplified version of) our new algorithm, together with our clustering and a generalization of match graphs. In Sections 5 and 6, we show how to reduce the space and implement our algorithm with a black-box finite automaton simulation. Finally, in Section 7, we show how to extend our results to efficiently support the interval operator.
@@ 2 Regular Expressions and Finite Automata
We review the classical concepts used in the paper. For more details, see e.g., Hopcroft and Ullman [30].
@@
The parse tree $T(R)$ of $R$ (not to be confused with the parse of $Q$ wrt. $R$) is the rooted, binary tree representing the hierarchical structure of $R$. The leaves of $T(R)$ are labeled by a character from $\Sigma\cup\{\epsilon\}$ and internal nodes are labeled by either $\odot$, $\mid$, $\cap$, $\neg$, or . For simplicity in our presentation, we identify extended regular expressions with their parse tree. We denote the subtree (or equivalently the subexpression) rooted at $v$ by $R(v)$. See Figure 1 for an example. Given two nodes $v$ and $w$ in $T(R)$, the lowest common ancestor, denoted $\mathsf{lca}(v,w)$, is the common ancestor of $v$ and $w$ with the largest depth.
@@
Given a (non-extended) regular expression $R$, we can construct an NFA accepting precisely the strings in $L(R)$ by several classical methods [41, 26, 58]. In particular, Thompson [58] gave the simple textbook construction shown in Figure 2. We will call an NFA constructed with these rules a Thompson NFA (TNFA). A TNFA $N(R)$ for $R$ has at most $2m$ states, at most $4m$ transitions, and can be computed in $O(m)$ time. We can compute a state-set transition on a single character in $O(m)$ time with a breadth-first search. Hence, it follows that we can solve (non-extended) regular expression matching on $Q$ in $O(nm)$ time and $O(m)$ space using the above state-set transition algorithm.
@@ 3 Dynamic Programming Algorithm
Let $R$ be an extended regular expression of length $m$ and let $Q$ be a string of length $n$. We review Hopcroft and Ullman’s dynamic programming algorithm [30] and show how to implement it using fast matrix multiplication in $O(n^{\omega}m)$ time and $O(n^{2}m/w+n+m)$ space.
@@ 3.2 Algorithm
Finally, we report that $R$ matches $Q$ iff the edge $(g_{1},g_{n+1})$ is in $G(r)$, where $r$ is the root of $T(R)$.
By induction on the parse tree, it follows that the algorithm correctly determines if $R$ matches $Q$. To implement the algorithm efficiently, we store each match graph $G(v)$ as a boolean $(n+1)\times(n+1)$ matrix, such that entry $(i,j)$ is $1$ iff there is an edge $(g_{i},g_{j})$ in $G(v)$. Consider the cases in the bottom-up traversal. We have that case 2.1 is a single boolean matrix multiplication of the matrices representing $G(v_{1})$ and $G(v_{2})$, and case 3.2 is the transitive closure of the matrix representing $G(v_{1})$, which can also be computed in the same time as a single matrix multiplication [21]. The remaining cases are straightforward to implement in $O(n^{2})$ time. Thus, we use at most $O(n^{\omega})$ time on each node in the parse tree, and hence the total time is $O(n^{\omega}m)$. Since each matrix uses $O(n^{2})$ bits, the total space is $O(n^{2}m/w+n+m)$.